OpenStudy (anonymous):

(16m^2 + 40m + 25) ÷ (4m + 5) = _____

OpenStudy (ash2326):

@bcheyennehaynes What's the square of (4m+5) \[(4m+5)^2=?\]

OpenStudy (anonymous):

=(4m+5)^2 / (4m+5) = 4m+5

OpenStudy (anonymous):

16m^2 + 40m + 25=(4m+5)^2

OpenStudy (anonymous):

factor out first then cancel out the 4m+5